Find the distance between the points (-2,3) and (1,-1).
Answer by jim_thompson5910(35256)

Use the distance formula


d = sqrt((x2-x1)^2+(y2-y1)^2)


Remember that the two generic points are (x1, y1) and (x2, y2)

So (x1, y1) = (-2, 3), which means x1=-2 and y1=3

and (x2, y2) = (1, -1), which means x2=1 and y2=-1


So in short, x1=-2, y1=3, x2=1, and y2=-1.

-------------------


d = sqrt((x2-x1)^2+(y2-y1)^2)


d = sqrt((1-(-2))^2+(-1-3)^2)


d = sqrt((1+2)^2+(-1-3)^2)


d = sqrt((3)^2+(-4)^2)


d = sqrt(9+16)


d = sqrt(25)


d = 5


So exact distance between the two points is 5 units.

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------